Tohban Report 2016-06-29
Tohban Reports | |
---|---|
Start Date: | 22 June 2016 |
End Date: | 29 June 2016 |
Tohban: | Mitsuo Oka |
Tohban email: | moka@ssl.berkeley.edu |
Next Tohban: | Jim McTiernan |
List all reports |
Solar Activity
Solar activity was very low during the week. Currently, there is no active region on disk, and no significant solar activity is predicted in the following days.
How many GOES flares occurred?
Flares above B, C, M, X class were 11 0 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 7 0 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 5 0 0 0
There were RHESSI flares/GOES flares 36 / 11 over the time range 21-Jun-16 28-Jun-16
Memory Management
The SSR has been emptying at the end of the passes set.
Spacecraft Status
Detectors 1, 3 and 8 are active and taking data. The cold plates 1 and 2 temperatures are 134.9 and 133.4, respectively. The cold tip/plate temperatures have increased slightly after turning on D1. The cryocooler power is at ~76.7W.
Data Gaps
There was no significant (>300s) data gap.
Detector issues
D1 was turned on, and it segmented at around 2860 V. It is now taking events. No threshold changes.
If the average of cold tip 1&2 temperatures reached 113, we will turn off D1.
Other notes
Spacecraft Management
Decimation | Normal/Vigorous |
HLAT Decimation | No Rear decimation, no front decimation |
Night time data (fronts) | full-time |
Night time data (rears) | full-time |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | Working as expected |
Detector problems? | See notes above. |